How long will it take money to double if it invested at the following rates:
(1) 5.9% compounded quarterly.
(2) 6.3% compounded quarterly.
Thanks!
SuperDuper

Answer by mananth(16949)

Principal P = 100
Amount= 200
years=n
compounded (t) 4
Rate = 5.9 0.01
Amount = P*((n+r)/n)^n

200 = 100 *( 1 + 0.01 )^ n*t
2 = *( 1 + 0.01 )^ n* 4
ln 2 = 4 n *ln 1.01
0.69 = 4 0.01 *n
11.83 = n
11.83 years
-------------------------
Principal P = 100
Amount= 200
years=n
compounded (t) 4
Rate = 6.3 0.02
Amount = P*((n+r)/n)^n

200 = 100 *( 1 + 0.02 )^ n*t
2 = *( 1 + 0.02 )^ n* 4
ln 2 = 4 n *ln 1.02
0.69 = 4 0.02 *n
11.09 = n